# Predicting the suitability of photocatalysts for water splitting using Koopmans spectral functionals: The case of TiO2 polymorphs

### v1

@@ IV.2 Band alignment
PBE $7.22$ $5.55$ $7.11$ $4.90$ $7.17$ $4.78$
HSE0600footnotetext: [33]11footnotemark: 1 $8.66$ $4.99$- -- -- --
HSE06 (QM/MM)111[36] $7.83$- $8.3$- -- --
G0W0@PBE11footnotemark: 1 $7.29$ $3.03$- -- -- --
G0W0@HSE0611footnotemark: 1 $7.92$ $3.19$- -- -- --
scQPGW@HSE0611footnotemark: 1 $8.77$ $3.59$- -- -- --
KI@PBE $8.00$ $4.71$ $8.19$ $4.39$ $8.03$ $4.00$
pKIPZ@PBE $7.83$ $4.44$ $8.00$ $4.10$ $7.84$ $3.71$
exp -- ZPR $8.5$ $4.97-5.00$- -- -- --
exp222[80, 99, 100] $8.20$ $5.14-5.17$ 7.96, 8.07, 8.20 -- -- --
ZPR333[101] $-0.3$ $0.17$- -- -- --
The ionization potentials and electron affinities of the $\mathrm{TiO_{2}}$ polymorphs, as calculated via KI and pKIPZ following the method described in Section II.3, are reported in Table 2 alongside values obtained from other higher levels of theory and experiment. Once again it is important to account for ZPR when comparing against experiment. However, ZPR shifts for the VBM and CBM individually have only been reported for the rutile polymorph, calculated using a generalized Fröhlich model [101]. This technique is more approximate than those that were used to calculate the ZPR shifts in the band gaps listed in Table 1. Note that the combining the shifts for the VBM and CBM predicted by the Fröhlich model results in a ZPR reduction of the band gap (0.47 eV) that is up to 50% larger than those listed in Table 1, so these results should be treated with caution.
That caveat aside, for rutile KI@PBE underestimates the shifted experimental IP by 0.5 eV — comparable to the accuracy of G0W0@HSE06. The EA, on the other hand, is predicted by KI@PBE within the range of values reported experimentally — a marked improvement upon the GW results, and matched only by HSE06.
For anatase and brookite, no ZPR results for the IP and EA were found. The KI and pKIPZ result for the IP of anatase are fractionally lower than that given by HSE06 [36].
@@
The band alignment diagram represented in Figure 4 shows the IPs and EAs of the three polymorphs compared against the two redox reaction potentials. The H+/H2 potential lies at -4.44 eV [102], while the H2O/O2 potential is 1.23 eV below this (as per Equation 1) [8]. The positions of the valence and conduction bands should straddle these two potentials i.e. the VBM and CBM should not lie within the two dashed lines in Figure 4. According to our calculations, rutile TiO2 appears not to satisfy this condition, which may explain why — despite having a more desirable band gap than anatase and brookite — this polymorph is a slightly worse catalyst, especially compared to anatase [103]. (Of course, an optimal band alignment does not guarantee photocatalytic success: anatase also exhibits higher surface activity and desirable excitonic properties [103, 104].) Brookite, on the other hand, shows some promise as a photocatalyst: the positions of the valence and the conduction bands seem to be favorable, but its fractionally larger band gap would limit the amount of photons that would be absorbed when subjected to sunlight. In this case, one might be able to engineer the band gap.
The KI and pKIPZ results are qualitatively the same, with pKIPZ reporting a slightly more favorable CBM for rutile relative to the hydrogen reduction reaction. Still, with respect to the value of band gap and band edge positions, both KI and pKIPZ predict that anatase appears to be the most promising photocatalyst of the three.
@@ VI Acknowledgments
MS acknowledges support from EXAF (Excellence in Africa) Research Centre as part of the collaboration ”Embedded exact quantum dynamics for photocatalytic water splitting”. This research was supported by the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 205602).

### v2

@@ IV.2 Band alignment
PBE $7.29$ $5.41$ $7.43$ $5.16$ $7.17$ $4.75$
HSE0600footnotetext: [79]11footnotemark: 1 $8.66$ $4.99$ - - - -
HSE06 (QM/MM)111[97] $7.83$ - $8.3$ - - -
G0W0@PBE11footnotemark: 1 $7.29$ $3.03$ - - - -
G0W0@HSE0611footnotemark: 1 $7.92$ $3.19$ - - - -
scQPGW@HSE0611footnotemark: 1 $8.77$ $3.59$ - - - -
KI@PBE $8.38$ $4.34$ $8.59$ $4.08$ $8.34$ $3.71$
pKIPZ@PBE $8.14$ $3.95$ $8.33$ $3.69$ $8.08$ $3.30$
exp - ZPR $8.5$ 4.975.004.97-5.00 - - - -
exp222[108, 56, 109] $8.20$ 5.145.175.14-5.17 7.96, 8.07, 8.20 - - -
ZPR333[26] 0.3-0.3 $0.17$ - - - -
The ionization potentials and electron affinities of the $\mathrm{TiO_{2}}$ polymorphs, as calculated via KI and pKIPZ following the method described in Section II.3, are reported in Table 2 alongside values obtained from other higher levels of theory and experiment. Once again it is important to account for ZPR when comparing against experiment. However, ZPR shifts for the VBM and CBM individually have only been reported for the rutile polymorph, calculated using a generalized Fröhlich model [26]. This technique is more approximate than those that were used to calculate the ZPR shifts in the band gaps listed in Table 1. Note that the combining the shifts for the VBM and CBM predicted by the Fröhlich model results in a ZPR reduction of the band gap (0.47 eV) that is up to 50% larger than those listed in Table 1, so these results should be treated with caution. That caveat aside, for rutile KI@PBE underestimates the shifted experimental IP by 0.5 eV — comparable to the accuracy of G0W0@HSE06. The EA, on the other hand, is predicted by KI@PBE within the range of values reported experimentally — a marked improvement upon the GW results, and matched only by HSE06. For anatase and brookite, no ZPR results for the IP and EA were found. The KI and pKIPZ result for the IP of anatase are fractionally lower than that given by HSE06 [97].
@@
The band alignment diagram represented in Figure 5 shows the IPs and EAs of the three polymorphs compared against the two redox reaction potentials. The H+/H2 potential lies at -4.44 eV [110], while the H2O/O2 potential is 1.23 eV below this (as per Equation 1) [96]. The positions of the valence and conduction bands should straddle these two potentials i.e. the VBM and CBM should not lie within the two dashed lines in Figure 5. According to our calculations, rutile TiO2 appears to satisfy this condition, but only marginally, as its conduction-band minimum lies only slightly above the reduction potential. This borderline alignment may help explain why—despite having a more desirable band gap than anatase and brookite—rutile often performs as a less efficient photocatalyst, particularly relative to anatase[70]. Brookite, on the other hand, shows some promise as a photocatalyst: the positions of the valence and the conduction bands seem to be favorable, but its fractionally larger band gap would limit the amount of photons that would be absorbed when subjected to sunlight. In this case, one might be able to engineer the band gap. The KI and pKIPZ results are qualitatively the same, with pKIPZ reporting a slightly more favorable CBM for rutile relative to the hydrogen reduction reaction. Still, with respect to the value of band gap and band edge positions, both KI and pKIPZ predict that anatase appears to be the most promising photocatalyst of the three.
Of course, an optimal band alignment does not guarantee photocatalytic success: anatase also exhibits higher surface activity and desirable excitonic properties [70, 31]. Indeed, excitonic binding energies and lifetimes vary significantly across the polymorphs. Rutile hosts weakly bound excitons (4–26 meV), with experiments reporting a 4 meV exciton [89] and GW calculations placing the lowest dipole-allowed exciton at 26 meV [55]; these small binding energies allow thermal dissociation, and charge carriers in rutile exhibit very short lifetimes [120]. In contrast, anatase supports more strongly bound excitons, with GW calculations predicting a 160 meV exciton in the (001) plane [6], and a recent study including lattice distortions reporting exciton–polaron binding energies of 216 meV [23]. Experiments also indicate that charge carriers in anatase are substantially longer-lived than in rutile [120]. (To our knowledge, excitonic data for brookite are not available.) Importantly, these excitonic effects are not strong enough to appreciably shift the band-edge positions relative to the redox potentials and therefore do not influence the band-alignment results presented here.
@@ VI Acknowledgments
MS acknowledges support from EXAF (Excellence in Africa) Research Centre as part of the collaboration “Embedded exact quantum dynamics for photocatalytic water splitting”. This research was supported by the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 205602).